Jason is 25 years old and has a new job in Web development. He wants to make sure he is financially sound in 30 years, so he plans to invest the same amount into a retirement account at the end of every month for the next 30 years. Suppose he wants to accumulate $5,000,000 when he retires. Create the spreadsheet model, then answer the question that how much he should invest at the end of each month to achieve the goal. Must show the spreadsheet model for any credit. Assume that the annual interest rate for the entire period is 10%.
Answer #1

Desired Future Value, FV = $5,000,000.00

Interest Rate, R % = 10

Remitting Frequency per year, N = 12

Time in Years, T = 30

Interest Rate per Period, ι = R/100N = 0.008333333

Number of Periods, p = NT = 360

PMT (Ordinary) = FV * ι / [{(1 + ι)^p} - 1] = $2,211.91

T LR/100N p NT PMT (Ordinary) FV/p-11 30 0.0083333 360 FV 5000000 10 12 2211.911838